# Offline-to-Online Multi-Agent Reinforcement Learning with Offline Value Function Memory and Sequential Exploration

### v1

@@ Abstract.
Multi-Agent Reinforcement Learning, Offline-to-Online Reinforcement Learning
@@ 1. Introduction
Multi-Agent Reinforcement Learning (MARL) has achieved remarkable success across various domains, including mastering complex video games Vinyals et al. (2019); Berner et al. (2019); Baker et al. (2020), optimizing warehouse logistics Shen et al. (2023), enabling robotic soccer Liu et al. (2022), and performing bi-manual dexterous manipulation Chen et al. (2024). However, these successes often come at the cost of low sample efficiency and high computational overhead, as MARL algorithms must explore a joint state-action space that grows exponentially with the number of agents. A promising approach to alleviate this computational burden is Offline-to-Online (O2O) Reinforcement Learning (RL). In recent years, O2O RL has achieved significant progress Zhang et al. (2024); LEI et al. (2024); Wang et al. (2023b); Hu et al. (2024b); Nakamoto et al. (2023); Zhao et al. (2024); Zhang et al. (2023); Lee et al. (2021), leveraging advances in Offline RL by utilizing offline datasets Fu et al. (2020); Formanek et al. (2023) to provide strong initial policies and pre-trained value functions. By learning a high-quality pre-trained policy from offline data, single-agent O2O RL can significantly reduce the need for extensive exploration and further enhance performance through interaction during the online phase. However, existing results primarily focus on single-agent scenarios, and the important O2O MARL setting has received only very limited attention, e.g., Jiang and Lu (2023).
O2O MARL faces two major challenges. First, exploration during the transition from the offline-to-online phase introduces a distributional shift, that can result in significant unlearning of the pre-trained Q-values in the early stages of online learning. To demonstrate this phenomenon, we conducted an experiment analyzing how pre-trained Q-values evolve during online fine-tuning. We collected test trajectories for the offline pre-trained agent interacting with the environment and stored them into a replay buffer. During the online phase, we evaluated how online value estimations for these samples evolved. As shown in Figure 1, Multi-agent Conservative Q-Learning (MACQL) Kumar et al. (2020), Multi-agent Cal-QL (MACal-QL) Nakamoto et al. (2023), and the action proposal method from Hu et al. (2024a) exhibit clear signs of unlearning, as the fine-tuned Q-values exhibit a rapid drop during the initial stage of online learning. This suggests that the algorithms have forgotten the previously learned optimal actions. This unlearning behavior hinders the efficiency of online fine-tuning, as the policy has to relearn knowledge that was already acquired during the offline phase. In contrast, our proposed algorithm OVMSE preserves the offline knowledge and achieves fast online fine-tuning.
@@
To address these two challenges, we propose a novel O2O MARL algorithm named Offline Value Function Memory and Sequential Exploration (OVMSE). Specifically, OVMSE consists of two key components. The first component, OVM, introduces a novel way of computing target Q-values and solves the unlearning problem. With OVM, even when the offline pre-trained Q-values degrade due to distributional shifts during the early stages of online learning, the algorithm allows for the quick recovery of the offline pre-trained values. The second key component, Sequential Exploration, is designed to enable efficient exploration after transitioning to the online phase. Inspired by the sequential update mechanism Bertsekas (2021); Wang et al. (2023a), SE restricts exploration to one agent at a time by allowing a single agent to select a random action while the others follow their respective policies. SE reduces the complexity of the joint state-action space and improves the quality of exploration, enabling more efficient fine-tuning and better performance. We further develop a decentralized SE for execution. We empirically validate OVMSE on a collection of easy, hard, and super-hard tasks in SMAC. Our results demonstrate superior performance compared to baseline methods.
The paper is organized as follows. Section 2 discusses related literature, Section 3 introduces the background knowledge for MARL and QMIX Rashid et al. (2018), Section 4 describes the proposed OVMSE algorithm, and Section 5 presents the experimental results and discussions. The main contributions of our paper are summarized as follows:
@@ 2. RELATED WORK
O2O RL and O2O MARL. Many previous works on O2O RL focus on reusing offline data during online training Ball et al. (2023); Lee et al. (2021); Hu et al. (2024b). Zhang et al. (2023) proposes retaining a copy of the offline policy and using a hybrid of the offline and online fine-tuned policies during online training, thereby preserving the performance achieved during the offline phase. A related approach to our work is presented in Hu et al. (2024a) which utilizes the offline policy to suggest actions for online target value estimation. In the context of O2O MARL, relatively little research attention has been given. Jiang and Lu (2023) investigates O2O MARL by using offline data during the online phase and proposes metrics for sampling from offline data. In addition to our proposed value function memory, we employ a multi-agent sequential exploration strategy inspired by multi-agent sequential rollout Bertsekas (2021); Wang et al. (2023a) to reduce the exploration space in cooperative multi-agent systems. This improves online learning efficiency, as the offline pre-trained policy can focus on targeted exploration rather than an exhaustive random search of the action space, which is typically required when training from scratch.
Offline MARL. The principles used to address accumulated extrapolation error in offline single-agent RL also apply to offline MARL. A common strategy is to incorporate pessimism towards out-of-distribution (OOD) states and actions, which has proven effective in the offline MARL setting as well. Both Pan et al. (2022) and Shao et al. (2023) extend the Conservative Q-Learning (CQL) framework Kumar et al. (2020) to the multi-agent domain. Pan et al. (2022) identifies that value functions in MARL are more prone to local optima as the number of agents increases and proposes using zeroth-order optimization to avoid being trapped in these local optima. Shao et al. (2023) demonstrates that independently penalizing each agent’s OOD actions can reduce the overall conservatism in value estimation, leading to improved empirical performance. Yang et al. (2021) takes a different approach by ensuring that only actions present in the dataset are used for target Q-value estimation, thereby promoting conservatism for OOD states and actions. Another promising direction is to treat offline MARL problems as supervised learning tasks. For example, Tseng et al. (2022) and Meng et al. (2023) frame offline MARL as sequence modeling problems, leveraging the power of decision transformers Chen et al. (2021) to achieve strong empirical performance. Furthermore, Li et al. (2023) and Zhu et al. (2023) employ diffusion models as policy backbones, demonstrating good performance and the ability to learn diverse strategies.
@@ 3.1. Multi-Agent Reinforcement Learning Formulation
A cooperative MARL problem can be described as a decentralized partially observable Markov decision process (Dec-POMDP) Oliehoek and Amato (2016), which can be represented as a tuple $\langle\mathcal{N},S,O,A,P,r,\gamma,T\rangle$.
@@ 3.2. QMIX
In this section, we present QMIX Rashid et al. (2018, 2020), which is a popular MARL algorithm following the Centralized Training with Decentralized Execution (CTDE) paradigm. In our work, we utilize QMIX as the backbone for both the offline and online phases.
@@
thereby maintaining that the joint action-value function $Q_{\text{tot}}$ is a non-decreasing function with respect to each individual action-value function $Q_{i}$. Hence, QMIX satisfies the Individual-Global-Maximum (IGM) property Son et al. (2019), which guarantees that the joint optimal action can be constructed by combining each agent’s individual optimal actions based on their local observations:
@@ 4.1. Offline Value Function Memory
In the online phase, we utilize QMIX as the backbone of our framework, replacing the Q target value with our Offline Value Function Memory (OVM) target. We describe the functionality of our proposed OVM training objective for online training. After offline training, we retain a copy of the pre-trained target value function, denoted as $\bar{Q}_{\text{tot-offline}}$. We introduce the OVM target, defined as:
@@
We further propose an annealing schedule for $\lambda_{\text{memory}}$, which gradually adjusts the influence of the offline value memory over time. The annealing schedule is defined as follows:
$\lambda_{\text{memory}}$ $=\min\Bigg{(}\lambda_{\text{memory\_end}},\,\left(1-\left(\frac{1-\lambda_{\text{memory\_end}}}{T}\right)\cdot t\right)\Bigg{)}$ (6)
@@
Figure 2. Median test win rates with respect to environment steps for OVMSE and baselines across different tasks pre-trained with medium or medium replay datasets. Our results demonstrate that OVMSE preserves offline knowledge, enables efficient exploration, and leads to faster fine-tuning and superior overall performance.
@@ 4.2. Multi-agent Sequential Exploration
$\epsilon$-greedy exploration is one of the most widely used strategies for exploration in multi-agent reinforcement learning (MARL) algorithms Hu et al. (2023); Rashid et al. (2018). In this strategy, each agent independently selects a random action with probability $\epsilon_{t}$, which gradually decreases over environment steps:
$\epsilon_{t}=\min\Bigg{(}\epsilon_{\text{end}},\,\left(1-\left(\frac{\epsilon_{\text{start}}-\epsilon_{\text{end}}}{T}\right)\cdot t\right)\Bigg{)},$ (7)
@@
Inspired by the multi-agent sequential update scheme Bertsekas (2021); Wang et al. (2023a), which sequentially updates different agents’ policies, we introduce the concept of sequential exploration for O2O MARL. Specifically, we adapt this concept into a sequential $\epsilon$-greedy exploration strategy. At each time step, agents collectively decide whether to explore based on the probability $\epsilon_{t}$, which decays with the number of environment steps. If exploration is chosen, only one randomly selected agent performs a random action, while all other agents act greedily, following the current policy. This coordinated approach allows the system to efficiently leverage the pre-trained policy while facilitating targeted exploration to further refine and improve the policy during the online phase.
Decentralized Sequential Exploration. However, this coordinated exploration requires communication between agents during execution, which may not always be feasible in decentralized execution settings. To address this limitation, we introduce a decentralized version of sequential exploration with a simple modification. We define a decentralized exploration probability, $\epsilon_{\text{dec\_t}}=\epsilon_{t}/N$, where $N$ is the number of agents. Each agent independently decides to explore with probability $\epsilon_{\text{dec\_t}}$, ensuring that, on average, the number of agents exploring at any given time matches that of the centralized version. This decentralized approach maintains the benefits of sequential exploration while ensuring compatibility with decentralized execution environments.
@@ 5.1. Setup
Codebase, Offline Datasets, and Experiment Setup. We evaluate our proposed OVMSE algorithm using the StarCraft Multi-Agent Challenge (SMAC), a key benchmark for MARL algorithms, across four tasks: 2s3z (easy), 3s5z (easy), 5m_vs_6m (hard), and 6h_vs_8z (super hard). For the tasks 2s3z, 5m_vs_6m, and 6h_vs_8z, we use offline datasets provided by CFCQL Shao et al. (2023). We collected our own dataset for 3s5z, as it is not publicly available. Offline training leverages medium and medium-replay datasets Fu et al. (2020). During the online phase, we apply the CQL loss for OVMSE in 2s3z, as it stabilizes online training. However, for 3s5z, 5m_vs_6m, and 6h_vs_8z, we exclude the CQL loss because it hinders online learning by preventing the discovery of new strategies. Our implementation, including code and hyperparameters, is based on PYMARL2 Hu et al. (2023) and CFCQL. Fine-tuning is conducted for 1 million environment steps in 2s3z and 3s5z, and for 3 million steps in 5m_vs_6m and 6h_vs_8z.
Baselines.We compare OVMSE against four baselines: (i) MACQL Kumar et al. (2020), which uses CQL loss throughout both the offline and online phases; (ii) MACal-QL, which applies Cal-QL (an algorithm designed for O2O single-agent RL) Nakamoto et al. (2023) to MARL; (iii) QMIX, trained from scratch Rashid et al. (2018); and (iv) Switch CQL in 3s5z, 5m_vs_6m, and 6h_vs_8z, which switches off the CQL loss during the online phase. All baselines use QMIX as the backbone.
Mixing Ratio and Hyperparameters. Following Nakamoto et al. (2023), we introduce the concept of a mixing ratio, a value ranging from 0 to 1, representing the percentage of offline data used during online training. For instance, a mixing ratio of 0 means no offline data is used during online training, while a ratio of 0.5 means half the data comes from offline and half from online sources. In this work, we select the optimal mixing ratio hyperparameters for all methods from the set $\{{0.0,0.1,0.3,0.5}\}$. Our experiments show that OVMSE performs best with a mixing ratio of 0.0 or 0.1. The detailed hyperparameter settings can be found in Appendix section A.
Results. We report the median test win rates over environment steps in Figure 2 and the mean test return in Figure 3. Additionally, we conduct 640 tests for each seed after online fine-tuning, with the final median test win rates presented in Table 1.
Table 1. The final median win rates for different tasks across algorithms, except for 2s3z in which all algorithms achieve near $100\%$ win rate. The results are obtained based on $640$ tests for each seed after the completion of online fine-tuning.
@@ 5.2. Results
Figure 2 and 3 present the training curves across four different tasks, pre-trained using both the medium and medium replay datasets. Our results show that OVMSE enables faster online fine-tuning and achieves superior overall performance. We compare OVMSE against baseline methods across three key aspects: (i) online performance, (ii) sample efficiency, and (iii) performance drop during the offline-to-online transition.
Table 1 shows the final fine-tuned performance for all tasks (except for 2s3z, where all algorithms achieve near $100\%$. win rates). In the 6h_vs_8z medium replay task, OVMSE outperforms QMIX and Switch-CQL by more than $20\%$. in win rate. Similarly, in the 3s5z medium tasks, OVMSE leads other baselines by roughly $10\%$ in win rate, and in the 3s5z medium replay, 6h_vs_8z medium, and 5m_vs_6m medium tasks, OVMSE demonstrates a lead of more than $5\%$ in test win rate.

### v2

@@ Abstract.
Multi-Agent Reinforcement Learning; Offline-to-Online Reinforcement Learning
[…]
@@ 1. Introduction
Multi-Agent Reinforcement Learning (MARL) has achieved remarkable success across various domains, including mastering complex video games grandmasterstarcraft2; Dota; EmergentToolUse, optimizing warehouse logistics logisticswarehouse, enabling robotic soccer deepmindscooer, and performing bi-manual dexterous manipulation Bidexhands. However, these successes often come at the cost of low sample efficiency and high computational overhead, as MARL algorithms must explore a joint state-action space that grows exponentially with the number of agents. A promising approach to alleviate this computational burden is Offline-to-Online (O2O) Reinforcement Learning (RL). In recent years, O2O RL has achieved significant progress Aperspectiveofqlearning; Unio4; PopulationO2O; BayesianO2O; Cal-QL; ENOTO; PEX; lee2021offlinetoonline, leveraging advances in Offline RL by utilizing offline datasets d4rl; offthegrid to provide strong initial policies and pre-trained value functions. By learning a high-quality pre-trained policy from offline data, single-agent O2O RL can significantly reduce the need for extensive exploration and further enhance performance through interaction during the online phase. However, existing results primarily focus on single-agent scenarios, and the important O2O MARL setting has received only very limited attention, e.g., MAO2O.
O2O MARL faces two major challenges. First, exploration during the transition from the offline-to-online phase introduces a distributional shift, that can result in significant unlearning of the pre-trained Q-values in the early stages of online learning. To demonstrate this phenomenon, we conducted an experiment analyzing how pre-trained Q-values evolve during online fine-tuning. We collected trajectories for the offline pre-trained agent interacting with the environment and stored them into a replay buffer. During the online phase, we evaluated how online value estimations for these samples evolved. As shown in Figure 1, Multi-agent Conservative Q-Learning (MACQL) CQL, Multi-agent Cal-QL (MACal-QL) Cal-QL, and the action proposal method from ImitationBoostrap exhibit clear signs of unlearning, as the fine-tuned Q-values exhibit a rapid drop during the initial stage of online learning. This suggests that the algorithms have forgotten the previously learned optimal actions. This unlearning behavior hinders the efficiency of online fine-tuning, as the policy has to relearn knowledge that was already acquired during the offline phase. In contrast, our proposed algorithm OVMSE preserves the offline knowledge and achieves fast online fine-tuning.
@@
To address these two challenges, we propose a novel O2O MARL algorithm named Offline Value Function Memory and Sequential Exploration (OVMSE). Specifically, OVMSE consists of two key components. The first component, OVM, introduces a novel way of computing target Q-values and solves the unlearning problem. With OVM, even when the offline pre-trained Q-values degrade due to distributional shifts during the early stages of online learning, the algorithm allows for the quick recovery of the offline pre-trained values. The second key component, Sequential Exploration, is designed to enable efficient exploration after transitioning to the online phase. Inspired by the sequential update mechanism MASEQROLLOut; OrderMatters, SE restricts exploration to one agent at a time by allowing a single agent to select a random action while the others follow their respective policies. SE reduces the complexity of the joint state-action space and improves the quality of exploration, enabling more efficient fine-tuning and better performance. We further develop decentralized SE for decentralized execution. We empirically validate OVMSE on a collection of easy, hard, and super-hard tasks in SMAC. Our results demonstrate superior performance compared to baseline methods.
The paper is organized as follows. Section 2 discusses related literature, Section 3 introduces the background knowledge for MARL and QMIX QMIX, Section 4 describes the proposed OVMSE algorithm, and Section 5 presents the experimental results and discussions. The main contributions of our paper are summarized as follows:
@@ 2. RELATED WORK
O2O RL and O2O MARL. Many previous works on O2O RL focus on reusing offline data during online training RLPD; lee2021offlinetoonline; BayesianO2O. PEX proposes retaining a copy of the offline policy and using a hybrid of the offline and online fine-tuned policies during online training, thereby preserving the performance achieved during the offline phase. A related approach to our work is presented in ImitationBoostrap which utilizes the offline policy to suggest actions for online target value estimation. In the context of O2O MARL, relatively little research attention has been given. MAO2O investigates O2O MARL by using offline data during the online phase and proposes metrics for sampling from offline data. In addition to our proposed value function memory, we employ a multi-agent sequential exploration strategy inspired by multi-agent sequential rollout MASEQROLLOut; OrderMatters to reduce the exploration space in cooperative multi-agent systems. This improves online learning efficiency, as the offline pre-trained policy can focus on targeted exploration rather than an exhaustive random search of the action space, which is typically required when training from scratch.
Offline MARL. The principles used to address accumulated extrapolation error in offline single-agent RL also apply to offline MARL. A common strategy is to incorporate pessimism towards out-of-distribution (OOD) states and actions, which has proven effective in the offline MARL setting as well. Both OMAR and CFCQL extend the Conservative Q-Learning (CQL) framework CQL to the multi-agent domain. OMAR identifies that value functions in MARL are more prone to local optima as the number of agents increases and proposes using zeroth-order optimization to avoid being trapped in these local optima. CFCQL demonstrates that independently penalizing each agent’s OOD actions can reduce the overall conservatism in value estimation, leading to improved empirical performance. MAICQ takes a different approach by ensuring that only actions present in the dataset are used for target Q-value estimation, thereby promoting conservatism for OOD states and actions. Another promising direction is to treat offline MARL problems as supervised learning tasks. For example, MAKD and MADT frame offline MARL as sequence modeling problems, leveraging the power of decision transformers DecisionTransformer to achieve strong empirical performance. Furthermore, BeyondConservatism and MADiff employ diffusion models as policy backbones, demonstrating good performance and the ability to learn diverse strategies.
@@ 3.1. Multi-Agent Reinforcement Learning Formulation
A cooperative MARL problem can be described as a decentralized partially observable Markov decision process (Dec-POMDP) IntroDecPomdp, which can be represented as a tuple $\langle\mathcal{N},S,O,A,P,r,\gamma,T\rangle$.
@@ 3.2. QMIX
In this section, we present QMIX QMIX; QMIXJMLR, which is a popular MARL algorithm following the Centralized Training with Decentralized Execution (CTDE) paradigm. In our work, we utilize QMIX as the backbone for both the offline and online phases.
@@
thereby maintaining that the joint action-value function $Q_{\text{tot}}$ is a non-decreasing function with respect to each individual action-value function $Q_{i}$. Hence, QMIX satisfies the Individual-Global-Maximum (IGM) property Qtran, which guarantees that the joint optimal action can be constructed by combining each agent’s individual optimal actions based on their local observations:
@@ 4.1. Offline Value Function Memory
In the online phase, we introduce our Offline Value Function Memory (OVM) target. We describe the functionality of our proposed OVM training objective for online training. After offline training, we retain a copy of the pre-trained target value function, denoted as $\bar{Q}_{\text{tot-offline}}$. We introduce the OVM target, defined as:
@@
We further propose an annealing schedule for $\lambda_{\text{memory}}$, which gradually adjusts the influence of the offline value memory over time. The annealing schedule is defined as follows:  [^1] We clarify that the annealing schedules in Eq. (6) and Eq. (7) utilize the $\max$ operator to ensure the parameters are lower-bounded by their terminal values, correcting a typo in an earlier version that used $\min$.
$\lambda_{\text{memory}}$ $=\max\Bigg(\lambda_{\text{memory\_end}},\,\left(1-\left(\frac{1-\lambda_{\text{memory\_end}}}{T}\right)\cdot t\right)\Bigg)$ (6)
@@
Figure 2. Median test win rates with respect to environment steps for OVMSE and baselines across different tasks, pre-trained with medium or medium replay datasets. Our results demonstrate that OVMSE preserves offline knowledge, enables efficient exploration, and leads to faster fine-tuning and superior overall performance.
@@ 4.2. Multi-agent Sequential Exploration
$\epsilon$-greedy exploration is one of the most widely used strategies for exploration in multi-agent reinforcement learning (MARL) algorithms hu2023rethinkingtheimplementation; QMIX. In this strategy, each agent independently selects a random action with probability $\epsilon_{t}$, which gradually decreases over environment steps:
$\epsilon_{t}=\max\Bigg(\epsilon_{\text{end}},\,\left(1-\left(\frac{\epsilon_{\text{start}}-\epsilon_{\text{end}}}{T}\right)\cdot t\right)\Bigg),$ (7)
@@
Inspired by the multi-agent sequential update scheme MASEQROLLOut; OrderMatters, which sequentially updates different agents’ policies, we introduce the concept of sequential exploration for O2O MARL. Specifically, we propose a sequential $\epsilon$-greedy exploration strategy. At each time step, agents collectively decide whether to explore based on the probability $\epsilon_{t}$, which decays with the number of environment steps. If exploration is chosen, only one randomly selected agent performs a random action, while all other agents act greedily, following the current policy. This coordinated approach allows the system to efficiently leverage the pre-trained policy while facilitating targeted exploration to further refine and improve the policy during the online phase.
Decentralized Sequential Exploration. However, this coordinated exploration requires communication between agents during execution, which may not always be feasible in decentralized execution settings. To address this limitation, we introduce a decentralized version of sequential exploration with a simple modification. We define a decentralized exploration probability, $\epsilon_{\text{dec\_t}}=\epsilon_{t}/N$, where $N$ is the number of agents. Each agent independently decides to explore with probability $\epsilon_{\text{dec\_t}}$, ensuring that, on average, the number of agents exploring at any given time matches that of the centralized version. This decentralized approach maintains the benefits of sequential exploration while ensuring compatibility with decentralized execution.
@@ 5.1. Setup
Codebase, Offline Datasets, and Experiment Setup. We evaluate our proposed OVMSE algorithm using the StarCraft Multi-Agent Challenge (SMAC), a key benchmark for MARL algorithms, across four tasks: 2s3z (easy), 3s5z (easy), 5m_vs_6m (hard), and 6h_vs_8z (super hard). For the tasks 2s3z, 5m_vs_6m, and 6h_vs_8z, we use offline datasets provided by CFCQL CFCQL. We collected our own dataset for 3s5z, as it is not publicly available. Offline training leverages medium and medium-replay datasets d4rl. During the online phase, we apply the CQL loss for OVMSE in 2s3z, as it stabilizes online training. However, for 3s5z, 5m_vs_6m and 6h_vs_8z, we exclude the CQL loss because it hinders online learning by preventing the discovery of new strategies. Our implementation, including code and hyperparameters, is based on PYMARL2 hu2023rethinkingtheimplementation and CFCQL. Fine-tuning is conducted for 1 million environment steps in 2s3z and 3s5z, and for 3 million steps in 5m_vs_6m and 6h_vs_8z.
Baselines.We compare OVMSE against four baselines: (i) MACQL CQL, which uses CQL loss throughout both the offline and online phases; (ii) MACal-QL, which applies Cal-QL (an algorithm designed for O2O single-agent RL) Cal-QL to MARL; (iii) QMIX, trained from scratch QMIX; and (iv) Switch CQL in 3s5z, 5m_vs_6m, and 6h_vs_8z, which switches off the CQL loss during the online phase. All baselines use QMIX as the backbone.
Mixing Ratio and Hyperparameters. Following Cal-QL, we introduce the concept of mixing ratio, a value ranging from 0 to 1, representing the percentage of offline data used during online training. For instance, a mixing ratio of 0 means no offline data is used during online training, while a ratio of 0.5 means half the data comes from offline and half from online interactions. In this work, we select the optimal mixing ratio hyperparameters for all methods from the set $\{{0.0,0.1,0.3,0.5}\}$. Our experiments show that OVMSE performs best with a mixing ratio of 0.0 or 0.1. The detailed hyperparameter settings can be found in Appendix.
Results. We report the median test win rates over environment steps in Figure 2 and the mean test return in Figure 3. All results are averaged across 6 random seeds. Additionally, we conduct 640 tests for each seed after online fine-tuning, with the final median test win rates presented in Table 1.
Table 1. The final median win rates for different tasks across algorithms, except for 2s3z in which all algorithms achieve near $100\%$ win rate. The results are obtained based on 640640 tests for each seed after the completion of online fine-tuning.
@@
Task Mixing Ratio 0.0 Mixing Ratio 0.1 Mixing Ratio 0.3
2s3z medium 19.58 19.64 18.41
Table 2. We conduct an ablation study for OVMSE with different mixing ratios on 2s3z medium and provide the mean test return at 200k environment steps. The results show that OVMSE has a minimal reliance on offline data (performing best with a mixing ratio of 0.0 and 0.1), as it preserves the offline learned Q-values.
@@ 5.2. Results
Figure 2 and 3 present the training curves across four different tasks, pre-trained using both the medium and medium replay datasets. Our results show that OVMSE enables fast online fine-tuning and achieves superior overall performance. We compare OVMSE against baseline methods across three key aspects: (i) online performance, (ii) sample efficiency, and (iii) performance drop during the offline-to-online transition.
Table 1 shows the final fine-tuned performance for all tasks (except for 2s3z, where all algorithms achieve near $100\%$ win rates). In the 6h_vs_8z medium replay task, OVMSE outperforms QMIX and Switch-CQL by more than $20\%$ in win rate. Similarly, in the 3s5z medium tasks, OVMSE leads other baselines by roughly $10\%$ in win rate, and in the 3s5z medium replay, 6h_vs_8z medium, and 5m_vs_6m medium tasks, OVMSE demonstrates a lead of more than $5\%$ in test win rate.